# Primordial Black Holes from Primordial Voids

### v1

@@ Abstract
Primordial black holes (PBHs) are a compelling dark matter candidate and a unique probe of small-scale cosmological fluctuations. Their formation is usually attributed to large positive curvature perturbations, which collapse upon Hubble re-entry during radiation domination. In this work we investigate instead the role of negative curvature perturbations, corresponding to the growth of primordial void (PV) like regions. Using numerical relativity simulations, we show that sufficiently deep PV can undergo a nonlinear rebounce at the center, generating an effective overdensity that eventually collapses into a PBH. We determine the critical threshold for this process for a variety of equations of state, and demonstrate that the resulting black holes obey a scaling relation analogous to the standard overdensity case. These results establish primordial voids as a novel channel for PBH formation and highlight their potential impact on PBH abundances and cosmological signatures.
@@ I Introduction
Primordial black holes (PBHs) Hawking (1971); Carr and Hawking (1974) are among the most promising candidates to explain the nature of dark matter, while also providing a unique probe of physics in the early Universe. Unlike astrophysical black holes, PBHs arise from non-stellar processes operating shortly after inflation, from processes that include the collapse of large adabiatic and iso-curvature perturbations  Shibata and Sasaki (1999); Passaglia and Sasaki (2022); Yoo et al. (2022), cosmic string loops Jenkins and Sakellariadou (2020); Helfer et al. (2019), or bubble collisions Jung and Okui (2024) (for reviews see Bagui et al. (2025); Escrivà et al. (2022); Carr and Kuhnel (2020)). Among all possibilities, the most widely studied and theoretically well-motivated scenario involves the collapse of curvature fluctuations during the radiation-dominated era Polnarev and Musco (2007); Escrivà et al. (2020, 2021), and other pressure(less) fluids Mack et al. (2007); Harada et al. (2016). These perturbations are typically seeded by quantum fluctuations during inflation, stretched to super-Hubble scales, which later re-enter the Hubble volume.
In this standard picture, attention has centered on positive curvature perturbations, corresponding to overdense regions of spacetime. For such profiles, the local density contrast grows non-linearly upon Hubble re-entry, and collapse occurs if a critical threshold is exceeded Carr (1975); Musco et al. (2005); Musco and Miller (2013); Escrivà et al. (2020, 2021). This framework underlies the bulk of PBH formation models, and besides being natural dark matter candidates, they can play an important role as seeds of supermassive black holes and be the sources of gravitational waves detected by the LIGO-Virgo-KAGRA experiments Sasaki et al. (2018); Carr and Kuhnel (2020); Bagui et al. (2025).
@@
In this work we carry out for the first time the studies on PBH formations from rebounces of PVs. We identify the critical amplitude of negative curvature fluctuations required for collapse, demonstrate that the resulting PBHs obey a scaling relation akin to the overdensity case, and map the competing outcomes of dispersal, shock formation, and black-hole collapse. These results open a new avenue for PBH phenomenology and highlight the need to include PVs when assessing PBH abundances and their observational consequences.
@@ II Primordial black holes from curvature fluctuations
Large curvature perturbations in the early Universe can arise from inflationary dynamics beyond the simplest slow-roll picture. In particular, non-attractor phases provide a natural way to amplify the curvature perturbation $\zeta$ on super-Hubble scales. Examples include ultra-slow roll inflation Figueroa et al. (2021); Pattison et al. (2021), constant-roll inflation Motohashi et al. (2020); Motohashi and Starobinsky (2017a, b); Tomberg (2023); Inui et al. (2024), or spectator field scenarios such as curvaton dynamics Pi and Sasaki (2023a). In these cases, the usual conservation of $\zeta$ outside the Hubble volume is broken, allowing its amplitude to grow and potentially reach the large values required for PBH formation. A useful organizing principle for these scenarios is provided by the so-called logarithmic duality Pi and Sasaki (2023b); Inui et al. (2025, 2024), which relates the non-Gaussian curvature perturbation to an underlying Gaussian field and connects various non-attractor mechanisms. Explicitly, the curvature takes the form
@@ II.1 Geometric classifications of fluctuations and dynamical collapse
$^{(3)}\,\mathrm{R}=-e^{-2\zeta(r)}\left(4\zeta^{\prime\prime}(r)+\frac{8}{r}\zeta^{\prime}(r)+2\left(\zeta^{\prime}(r)\right)^{2}\right).$ (5)
This classification is relevant for PBH formation, because a closed core would facilitate the collapse, leading to a lower critical amplitude threshold, whereas an open core would oppose it. In this light, positive and negative values of $\mu$ in Eq.(3) correspond to open and closed core-types, respectively, and therefore a higher collapse threshold is expected for the PVs sector.
@@ II.3 Compaction function and threshold criterion
An alternative diagnostic for identifying the formation of a black hole is the compaction function, which can be expressed as
@@
$C_{\ell}(r)=2\phi(w)\,r\,\zeta^{\prime}(r)\,,\quad\phi(w)\equiv\frac{(1+3w)}{5+3w}\,$, (19)
@@ III General Relativistic Hydrodynamics
Figure 3: Evolution of the density contrast $\rho_{E}/\rho_{\rm bkg})$ for negative curvature fluctuations re-entering the Hubble horizon during radiation domination $(\omega=1/3)$, obtained from numerical relativity simulations. Top panels: initial fluctuation with amplitude $\mu=-1.8$, which forms a PV that rebounds into a central overdensity and collapses into a PBH. Bottom panels: initial fluctuation with amplitude $\mu=-0.8$, which also forms a PV but whose rebound dissipates into sound waves. Underdensities and overdensities are represented by a blue-to-red colormap. The green circle denotes the formation of an apparent horizon.
@@ IV Numerical results
Our numerical simulations begins from a initial equal-time hypersurface representing a super-Hubble curvature fluctuation $\zeta(t,r)$ centered at the origin of our radial coordinate, embedded in a Friedmann–Lemaître–Robertson–Walker (FLRW) expanding background. The overall scale factor is initially set to one, $\mathfrak{a}(t_{0})=1$ with an expanding Hubble rate $H(t)=\frac{2}{3(1+\omega)}\frac{1}{t}$, where $t_{0}$ is an arbitrary initial time in code units. The characteristic scale of $\zeta(t_{0},r)$ is set to $k_{\star}^{-1}=100\,H^{-1}(t_{0})$, ensuring that the perturbation is well outside the Hubble horizon initially.
@@ IV.1 Dynamics of gravitational collapse
Figure 4: Mass scaling of PBH formation for negative (left pannel) and positive (right pannel) curvature fluctuations.
@@
For a positive amplitude $\mu$, the perturbation corresponds to an initial overdensity. While outside the Hubble horizon, the perturbation remains quasi-linear. Its subsequent evolution after horizon entry is characterized by rapid growth and strong self-gravity. The final outcome is then determined by the competition between this self-gravity and pressure gradients. For sufficiently large amplitude $\mu>\mu_{c}^{(+)}$, the gravity dominates and the region collapses to form a PBH, whereas when not, the pressure support is sufficient to prevent the collapse and the overdensity disperses into sound waves without forming a black hole.
In contrast, a negative amplitude $\mu$ corresponds to an initial underdensity. The central underdensity deepens progressively, reaching its maximum depth at Hubble crossing—a stage we refer to as a PV. Owing to the oscillatory structure of the sinc profile, a PV is always accompanied by a surrounding overdense shell. During the subsequent sub-Hubble evolution, this shell contracts under its self-gravity and the background pressure, driving matter inwards and causing the central region to rebound. In this scenario, pressure initially drives the contraction, in direct contrast to the positive $\mu$ case where it always opposes collapse. For a sufficiently negative amplitude $\mu<\mu_{c}^{(-)}$, the rebound is sufficiently strong to collapse into a PBH; otherwise, it disperses, generating sound waves. This mechanism highlights the qualitative difference between PBH formation from overdensities and PVs. Figure 3 illustrates the dynamics of PVs in both collapsing and dissipating scenarios, showing the energy density ratio $\rho_{E}/\rho_{\rm bkg}$ for times before Hubble crossing, at rebound, and during the final collapse or dissipation of the central feature.
@@ IV.2 Equation of state dependence and thresholds
The formation dynamics exhibit a strong dependence on the fluid pressure, parameterized by the equation of state $\omega$. We have computed the critical thresholds for PBH formation (see Table 1) from both overdensities and PVs across a range of $\omega$, from the nearly pressureless limit ($\omega\simeq 0$) to the stiff fluid limit ($\omega=1$). Figure 1 shows the numerical collapse thresholds $\mu_{\rm c}$ for representative $\omega$ values. As expected, the thresholds for PVs exceed those for overdensities, as gravitational collapse must overcome the expanding void geometry. Furthermore, we find that while positive-$\mu$ perturbations can form PBHs directly from Type I fluctuations, negative-$\mu$ perturbations require Type II fluctuations. All PBHs produced in the vicinity after $\left|\mu\right|>\left|\mu_{c}^{(\pm)}\right|$ are of Type A.
Furthermore, the thresholds for positive $\mu$ increase monotonically with $\omega$, elucidating the role of pressure to prevent gravitational collapse. Instead, thresholds in the PV formation channel display a non-monotonic behavior, peaking near the radiation-dominated case $\omega\simeq 1/3$ and decreasing for both softer and stiffer equations of state. We interpret this as a reflection of competing effects of pressure gradients and sound speed in the collapse dynamics. For $\omega<1/3$, the reduced fluid’s pressure facilitates the collapse once the rebounce is produced. Conversely, for $\omega>1/3$, the increased sound speed associated with higher pressure enables a faster compression rate, producing a more violent rebound that effectively facilitates PBH formation, even though pressure subsequently opposes collapse once the overdense core is generated.
@@
After comparison of the numerical thresholds with those predicted by the analytical prescription (22), our numerical results confirm a good agreement for the positive-curvature fluctuations scenario in the radiation domination case, with deviations under 5%. For negative-curvature fluctuations, however, we observe that the presciption predicts a lower threshold ($\mu_{c}^{\rm th}\sim-1.1$), inducing approximately a 20% error. This disagreement is not surprising as these analytical approaches do not capture the specific open-core geometry associated to PVs.
IV.3 Scaling of mass functions
$\frac{M_{\rm PBH}}{M_{\times}}=K\left|\mu-\mu_{c}\right|^{\gamma}~,$ (29)
@@ V Conclusions and Discussions
This work represents one of the first investigations into the previously overlooked phenomenology of underdensities that may have originated during cosmic inflation. Several interesting extensions remain for future work, including studying the effects of non-Gaussianities on the shape profile—particularly those that produce a negative skewness in the probability distribution function of $\zeta$, as these would enhance the abundance of PVs and their imprints in cosmological observations.
In addition, the question of whether PVs can form PBHs of type-B (from a ”separate universe”) has been left unanswered. Would these objects exist, they would correspond (at least initially) to a trapped region with an open-core geometry. It would be interesting to better undertenant the fate of such structures.
Because underdensities or voids are necessarily present in any model that sufficiently enhances the scalar powerspectrum, it is important to take into account their associated phenomenology. Some examples include the generation of sound waves and related gravitational wave signatures Ning et al. (2025); Zeng et al. (2025), the potentially relevant effects to matter-dominated epochs such for PBH reheating models Allahverdi et al. (2025), or even their potential role as catalysts for baryogenesis Baumann et al. (2007); Carr et al. (2021); García-Bellido et al. (2021); Despontin et al. (2024).
Finally, extending this analysis to scalar-field-dominated cosmologies would also be relevant for scenarios where PBHs form shortly after inflation, either during a phase of kinationCheng et al. (2025), or (p)reheating Khlopov et al. (1985); Martin et al. (2020); Padilla et al. (2025); Milligan et al. (2025).

### v2

@@ Abstract
Primordial black holes (PBHs) are a compelling dark matter candidate and a unique probe of small-scale cosmological fluctuations. Their formation is usually attributed to large positive curvature perturbations, which collapse upon Hubble re-entry during radiation domination. In this work we investigate instead the role of negative curvature perturbations, corresponding to the growth of primordial void (PV) like regions. Using numerical relativity simulations, we show that sufficiently deep PV can undergo a nonlinear rebounce at the center, generating an effective overdensity that eventually collapses into a PBH. We determine the critical threshold for this process for a variety of equations of state, and demonstrate that the resulting black holes masses obey a scaling relation analogous to the standard overdensity case. These results establish primordial voids as a novel channel for PBH formation and highlight their potential impact on PBH abundances and cosmological signatures.
@@ I Introduction
Primordial black holes (PBHs) Hawking (1971); Carr and Hawking (1974) are among the most promising candidates to explain the nature of dark matter, while also providing a unique probe of physics in the early Universe. Unlike astrophysical black holes, PBHs arise from non-stellar processes operating shortly after inflation, from processes that include the collapse of large adiabatic and iso-curvature perturbations  Shibata and Sasaki (1999); Passaglia and Sasaki (2022); Yoo et al. (2022), cosmic string loops Jenkins and Sakellariadou (2020); Helfer et al. (2019), or bubble collisions Jung and Okui (2024) (for reviews see Bagui et al. (2025); Escrivà et al. (2022); Carr and Kuhnel (2020)). Among all possibilities, the most widely studied and theoretically well-motivated scenario involves the collapse of curvature fluctuations during the radiation-dominated era Polnarev and Musco (2007); Escrivà et al. (2020, 2021), other pressure(less) fluids Mack et al. (2007); Harada et al. (2016), and scalar fields de Jong et al. (2022, 2023); Cheng et al. (2025); Milligan et al. (2025); Padilla et al. (2025). These perturbations are typically seeded by quantum fluctuations during inflation, stretched to super-Hubble scales, which later re-enter the Hubble volume.
In this standard picture, attention has centered on positive curvature perturbations, corresponding to overdense regions of spacetime. For such profiles, the local density contrast grows non-linearly upon Hubble re-entry, and collapse occurs if a critical threshold is exceeded Carr (1975); Musco et al. (2005); Musco and Miller (2013); Escrivà et al. (2020, 2021). This specific mechanism underlies the bulk of PBH formation models, and the potential observational consequences of PBHs, such as their role as dark matter, seeds for supermassive black holes, or sources for gravitational waves Sasaki et al. (2018); Carr and Kuhnel (2020); Bagui et al. (2025), are therefore predominantly assessed within this framework.
@@
In this work we carry out a dedicated study on PBH formations from rebounces of PVs. We identify the critical amplitude of negative curvature fluctuations required for collapse, demonstrate that the resulting PBHs masses obey a scaling relation akin to the overdensity case, and map the competing outcomes of dispersal, shock formation, and black-hole collapse. Extending on Germani and Montellà (2025), these results open a new avenue for PBH phenomenology and highlight the need to include PVs when assessing PBH abundances and their observational consequences.
@@ II Primordial black holes from curvature fluctuations
Large curvature perturbations in the early Universe can arise from inflationary dynamics beyond the simplest slow-roll picture. In particular, non-attractor phases provide a natural way to amplify the curvature perturbation $\zeta$ on super-Hubble scales. Examples include ultra-slow roll inflation Germani and Prokopec (2017); Figueroa et al. (2021); Pattison et al. (2021), constant-roll inflation Motohashi et al. (2020); Motohashi and Starobinsky (2017a, b); Tomberg (2023); Inui et al. (2024), or spectator field scenarios such as curvaton dynamics Pi and Sasaki (2023a). In these cases, the usual conservation of $\zeta$ outside the Hubble volume is broken, allowing its amplitude to grow and potentially reach the large values required for PBH formation. A useful organizing principle for these scenarios is provided by the so-called logarithmic duality Atal et al. (2019); Pi and Sasaki (2023b); Inui et al. (2025, 2024), which relates the non-Gaussian curvature perturbation to an underlying Gaussian field and connects various non-attractor mechanisms. Explicitly, the curvature takes the form
@@ II.1 Geometric classifications of fluctuations and dynamical collapse
$^{(3)}R=-e^{-2\zeta(r)}\left(4\zeta^{\prime\prime}(r)+\frac{8}{r}\zeta^{\prime}(r)+2\left(\zeta^{\prime}(r)\right)^{2}\right).$ (5)
This classification is relevant for PBH formation, because a closed core would facilitate the collapse, leading to a lower critical amplitude threshold, whereas an open core would oppose it. In this light, positive and negative values of $\mu$ in Eq.(3) correspond to closed and open core-types, respectively, and therefore a higher collapse threshold is expected for the PVs sector.
@@ II.3 Compaction function and threshold criterion
An alternative diagnostic for identifying the formation of a black hole is the compaction function Shibata and Sasaki (1999), which can be expressed as
@@
$C_{\ell}(r)=-2\phi(w)\,r\,\zeta^{\prime}(r)\,,\quad\phi(w)\equiv\frac{3(1+w)}{5+3w}\,$, (19)
@@ III General Relativistic Hydrodynamics
Figure 3: Evolution of the density contrast $\rho_{E}/\rho_{\rm bkg}$(upper rows), relative recession velocity with respect the background (middle rows) and normalized Ricci Scalar by a factor $\mathfrak{a}(t)^{2}$ (bottom rows) for negative curvature fluctuations re-entering the Hubble horizon during radiation domination $(\omega=1/3)$, obtained from numerical relativity simulations. Top panels: initial fluctuation with amplitude $\mu=-1.8$ and $\omega=1/3$, which forms a PV that rebounds into a central overdensity and collapses into a PBH. Bottom panels: initial fluctuation with amplitude $\mu=-0.8$ and $\omega=1/3$, which also forms a PV but whose rebound dissipates into sound waves. Underdensities are represented by a blue or purple colormaps, while overdensities by red or green colormaps. The yellow circle denotes the formation of an apparent horizon.
@@ IV Numerical results
Our numerical simulations begin on an initial equal-time hypersurface containing a super-Hubble curvature fluctuation $\zeta(t,r)$ centered at the origin of the radial coordinate, embedded in a Friedmann–Lemaître–Robertson–Walker (FLRW) expanding background. The background scale factor is $\mathfrak{a}(t)=\mathfrak{a}_{0}(t/t_{0})^{\sigma}$}, with Hubble rate $H_{\rm bkg}(t)=\sigma/t$, where $\sigma=2/[3(1+\omega)]$} for a perfect fluid of equation-of-state parameter $\omega$. For convenience, we set $t_{0}=1$ and $\mathfrak{a}_{0}=\mathfrak{a}(t_{0})=1$.
The characteristic scale of the initial curvature fluctuation is fixed by $k_{\star}^{-1}=100\,H_{\rm bkg}^{-1}(t_{0})$, ensuring that the perturbation is well outside the Hubble radius at the start of the simulation. Thus, we define the Hubble-crossing time $t_{H}$ as the instant when $\mathfrak{a}(t)H_{\rm bkg}(t)/k_{\star}=1$.
@@ IV.1 Dynamics of gravitational collapse
Figure 4: Mass scaling of PBH formation for negative (left panel) and positive (right panel) curvature fluctuations.
@@
For a positive amplitude $\mu$, the perturbation corresponds to an initial overdensity. While outside the Hubble horizon, the perturbation remains quasi-linear, i.e. following linear perturbation theory until $k_{\star}/\mathfrak{a}H_{\rm bkg}\sim 0.1$. Its subsequent evolution after horizon entry is characterized by rapid growth and strong self-gravity. The final outcome is then determined by the competition between this self-gravity and pressure gradients. For sufficiently large amplitude $\mu>\mu_{c}^{(+)}$, the gravity dominates and the region collapses to form a PBH, whereas when not, the pressure support is sufficient to prevent the collapse and the overdensity disperses into sound waves without forming a black hole.
In contrast, a negative amplitude $\mu$ corresponds to an initial underdensity. The central underdensity deepens progressively already on super-Hubble scales, where the negative curvature perturbation causes the inner region to expand faster than the outer shells, reaching its maximum depth at Hubble crossing—a stage we refer to as a PV. Owing to the oscillatory structure of the sinc profile, a PV is always accompanied by a surrounding overdense shell. During the subsequent sub-Hubble evolution, this shell contracts under its self-gravity and the background pressure, driving matter inwards and causing the central region to rebound. In this scenario, pressure initially drives the contraction, in direct contrast to the positive $\mu$ case where it always opposes collapse. For a sufficiently negative amplitude $\mu<\mu_{c}^{(-)}$, the rebound is sufficiently strong to collapse into a PBH; otherwise, it disperses, generating sound waves. This mechanism highlights the qualitative difference between PBH formation from overdensities and PVs. Figure 3 illustrates the dynamics of PVs in both collapsing and dissipating scenarios. The top rows show the energy density ratio $\rho_{E}/\rho_{\rm bkg}$ for times before Hubble crossing, at rebound, and during the final collapse or dissipation of the central feature. The middle rows show the relative recession velocity from the origin, i.e. $V_{\rm rec}=v^{r}+HR$, where H=K/3H=-K/3 is the local expansion rate, with respect to the background recession $V_{\rm rec,\,bkg}=H_{\rm bkg}R$. Finally, in the bottom rows, we also show the evolution of the Ricci curvature, i.e. $^{(3)}R\cdot\mathfrak{a}(t)^{2}$.
@@ IV.2 Equation of state dependence and thresholds
The formation dynamics exhibit a strong dependence on the fluid pressure, parameterized by the equation of state $\omega$. We have computed the critical thresholds for PBH formation (see Table 1) from both overdensities and PVs across a range of $\omega$, from the nearly pressureless limit ($\omega\simeq 0$) to the stiff fluid limit ($\omega=1$). Figure 1 shows the numerical collapse thresholds $\mu_{\rm c}$ for representative $\omega$ values. As expected, the thresholds for PVs exceed those for overdensities, as gravitational collapse must overcome the expanding void geometry. Moreover, we find that while positive-$\mu$ perturbations can form PBHs directly from Type I fluctuations, negative-$\mu$ perturbations require Type II fluctuations. All PBHs produced in the vicinity after $\left|\mu\right|>\left|\mu_{c}^{(\pm)}\right|$ are of Type A.
Furthermore, the thresholds for positive $\mu$ increase monotonically with $\omega$, elucidating the role of pressure to prevent gravitational collapse. Instead, thresholds in the PV formation channel display a non-monotonic behavior, peaking near the radiation-dominated case $\omega\simeq 1/3$ and decreasing for both softer and stiffer equations of state. We interpret this as a reflection of competing effects of pressure gradients and sound speed in the collapse dynamics. For $\omega<1/3$, the reduced fluid’s pressure facilitates the collapse once the rebounce is produced. Conversely, for $\omega>1/3$, the increased sound speed associated with higher pressure, i.e. $c_{s}\approx~\sqrt{\omega}$, enables a faster compression rate, producing a more violent rebound that effectively facilitates PBH formation, even though pressure subsequently opposes collapse once the overdense core is generated.
@@
After comparison of the numerical thresholds with those predicted by the analytical prescription (22), our numerical results confirm a good agreement for the positive-curvature fluctuations scenario in the radiation domination case, with deviations under 5%. For negative-curvature fluctuations, however, we observe that the prescription predicts a lower threshold [^1] We use the prescription for type-II fluctuations from 22, of which $\delta_{c}=a\kappa_{m}^{b}$} with a=0.51962{a=0.51962} and b=0.26687{b=0.26687}. ($\mu_{c}^{\rm th}\sim-1.1$), inducing approximately a 20% error. This disagreement is not surprising as these analytical approaches do not capture the specific open-core geometry associated to PVs.
@@
IV.3 Scaling of mass functions
$\frac{M_{\rm PBH}}{M_{\times}}=K\left|\mu-\mu_{c}\right|^{\gamma}~,$ (29)
@@ V Conclusions and Discussions
This work represents one of the first explorations of the phenomenology associated with negative curvature perturbations generated during inflation. An important direction for future research is a more systematic study of the shape of the curvature profile. The profile can be modified by the width of the primordial power spectrum as well as by local non-Gaussianities, both of which influence the relative depth of the core and the height of the surrounding ridge. For sufficiently broad spectra, non-spherical features may also become relevant, and the statistics tend to favour sharper, shell–dominated configurations. A deeper understanding of how these shape variations affect the compaction function and the role of the three-Ricci curvature in the subsequent dynamics should ultimately improve analytical prescriptions for collapse thresholds for both positive and negative curvature profiles. We leave these extensions for future work.
A related open question concerns the possible existence of type-B PBHs sourced by PVs, corresponding to a “separate universe” configuration. If such objects can form, they would initially appear as trapped regions with an open-core geometry, and it would be important to clarify whether these evolve toward stable PBHs or instead disperse. A dedicated study would be required to address this scenario.
Furthermore, because underdensities or voids are necessarily present in any model that sufficiently enhances the scalar power spectrum, it is important to account for their associated phenomenology. Relevant examples include the generation of sound waves and the resulting gravitational wave signatures Ning et al. (2025); Zeng et al. (2025), potential effects during matter-dominated epochs such as those appearing in PBH reheating models Allahverdi et al. (2025), or even their possible role in producing non-thermalized regions that act as catalysts for baryogenesis Baumann et al. (2007); Carr et al. (2021); García-Bellido et al. (2021); Despontin et al. (2024). Understanding the signatures associated with relatively large positive and negative curvature fluctuations may help not only to constrain inflationary models at low frequencies, but also to elucidate potential mechanisms that address outstanding cosmological puzzles.
Finally, it would be valuable to extend the present analysis to scalar-field dominated epochs. PBH formation shortly after inflation, whether during a phase of kination Cheng et al. (2025) or during (p)reheating Khlopov et al. (1985); Martin et al. (2020); Padilla et al. (2025); Milligan et al. (2025), involves dynamics that may differ significantly from those of a perfect fluid. Incorporating the effects of coherent scalar-field oscillations, scalar field gradients and nontrivial sound speeds would provide further insight on the viability of PBH formation (and other relics) during these early epochs.